# Generative AI and Firm Productivity:

### v2

@@ 2.1 The Economic Impact of Generative AI
Recent advances in GenAI have attracted considerable attention for their potential economic and social implications. A growing body of research documents GenAI’s ability to enhance individual productivity in simple and well-defined tasks, including mid-level writing [54], software development [56, 24], marketing copy generation [26], and legal analysis [22]. However, this literature largely focuses on individual-level impacts in laboratory settings, with limited empirical evidence of measurable productivity gains at the aggregate or firm level [1, 18]. In fact, the productivity dynamics observed in real-world firm adoption are likely more complex than those captured in lab environments due to technical, organizational, and market factors ([15]; [50]).
More importantly, most prior studies assess the productivity potential of GenAI from a supply-side perspective, emphasizing labor savings or improvements in worker efficiency, typically measured by decreases in average task completion time or increases in the average number of completed tasks [54, 26, 56, 41, e.g.,]. Even the limited studies employing real-world firm data, such as [15] and [50], largely rely on similar supply-side metrics. By contrast, very few studies have examined GenAI’s productivity effects through demand-side value creation, such as enhanced consumer experience and increased purchases.777[21], [31], [46], and [40] show that AI-generated ad copies or images in digital advertising can raise the number of clicks or click-through rates, but they provide no evidence on actual purchase behavior.
Furthermore, another important question concerns the asymmetric effects of GenAI across different user groups. Earlier waves of technological change, such as the adoption of computers and the Internet, were often described as “skill-biased” ([38]), disproportionately favoring skilled users while leaving unskilled workers behind ([12]; [3]; [6]; [2]). As of now, the heterogeneous impacts of GenAI on worker performance appear more context-dependent. On one hand, by lowering skill barriers, GenAI can promote inclusivity ([52]; [30]; [23]), as it benefits users with lower levels of skills and expertise ([15]; [54]; [56]; [43]). On the other hand, evidence from other studies points to the opposite outcome ([57]; [55]).
@@ 2.2 Friction Reduction in Online Marketplaces
A key friction in online marketplaces is asymmetric information: buyers often cannot directly verify product quality or seller reliability prior to purchase [45, 62]. To mitigate this challenge and facilitate consumer decision-making, platforms have traditionally relied on reputation and review systems to generate quality signals [17, 29, 32, 64]. However, these feedback mechanisms suffer from well-documented limitations, including grade inflation [53, 68], “cold start” problems [4], and score manipulation [49, 48]. Modern platforms are increasingly adopting advanced technologies and innovative market designs. For instance, AI models based on natural language processing are used to extract insights from textual reviews or to filter feedback for relevance, enabling more accurate inferences about product quality and consumer satisfaction [51, 47]. Moreover, curated provision of off-site social media information during consumer search has been shown to assist consumer decision-making and enhance platform revenues [35].
Consumers on digital platforms also face substantial frictions during online search, arising either from search costs, the effort and resources required to locate information, or search targetability, the effectiveness of search engines in retrieving the most relevant products. Such frictions can significantly influence market outcomes and market structures in online marketplaces [36, 42, 65, 16, 14, 5]. To mitigate search frictions, digital platforms have invested heavily in technology and design innovations aimed at encouraging consumer search behavior and improving match value. Prior research highlights several such advances in online search, including ranking algorithms [28, 63, 66], refinement tools like sorting and filtering [20, 34], machine-learning-driven personalized search [67], category-refinement-based precision improvements [69], and optimal search engine information layout [39].
In online marketplaces, advances in personalization and targeting technologies can also help reduce frictions by delivering products or content to users most likely to be interested. [8] develop theoretical models showing that improving advertisers’ targeting ability increases the number of consumer-product matches, thus enhancing the overall social value of advertising. Empirically, targeted advertising has been found more effective than untargeted approaches. For instance, [37] observe that display ads matched to website context significantly raise purchase intent, while [9] suggest that targeted ads are particularly valuable in lowering search friction when consumers would otherwise have difficulty discovering or learning about products. Extending these insights to recommendation systems, [60] demonstrate that removing personalized recommendations discourages consumer engagement and purchasing, especially for small sellers and niche consumers.
@@
3 Study Setting, Experimental Design and Data
@@ 3.2 Theoretical Framework
We model the impact of GenAI adoption on firm productivity through the lens of the standard Solow growth model [58]. Assume that output is produced according to a Cobb–Douglas production function
@@
In Section 3.3, we explain why these assumptions are likely to hold in our context. Under such conditions, the potential gains from GenAI adoption can be interpreted as a pure productivity shock, represented by an upward shift in the $A$ term of the production function. This interpretation is consistent with treatments of past general-purpose technologies (e.g., electrification or the internet), where adoption translated into improvements in TFP rather than capital accumulation. [^8] A natural question is whether GenAI adoption should be viewed as capital deepening rather than TFP growth. While, in principle, new investments in IT infrastructure (e.g., servers, GPUs, or proprietary model development) could expand the capital stock, in our context the platform already possessed the required infrastructure, and GenAI applications constituted incremental software upgrades operating on existing systems. Evidence from consumer-facing sectors further suggests that quality improvements—including reductions in information asymmetry, better matching, enhanced personalization and targeting, and product innovation—can raise revenue-based productivity without corresponding shifts in input units or costs [61, 13, 25].
@@ 3.3 Empirical Framework
GenAI-generated descriptions on top of those created by humans Yes
@@
Pre-sale Service Chatbot:
The experiment, conducted over a two-month period from September to October 2023, included a random sample of 44 thousand consumers who initiated pre-sale customer service requests for the platform’s self-sold products during the experimental period. These consumers were randomly divided into treatment and control groups. Consumers in the control group received the platform’s automated response service, which delivered a pre-programmed standardized notification indicating that customer service was unavailable. This auto-response condition reflects the platform’s standard operating practice. Due to resource considerations, the platform has historically prioritized allocating human agents to post-sale rather than pre-sale support for self-sold products, given that pre-sale inquiries are generally less urgent. By contrast, consumers in the treatment group were supported purely by GenAI-powered chatbots. GenAI is expected to reduce asymmetric information between buyers and sellers in the treatment group by providing richer, context-specific responses to consumer inquiries.
Search Query Refinement:
@@
The experiment comprised five sub-experiments, each involving consumers who spoke English, Spanish, French, Portuguese, or Korean, which are among the most widely used languages on the platform. All sub-experiments ran for one week in December 2023, with staggered start dates. GenAI was employed to create multilingual product descriptions for a predetermined product set of approximately 45,000 randomly selected platform self-sold products spanning a broad range of categories. On this platform, product descriptions refer to the text content in the description module of product detail pages that summarizes key features and selling points. According to our partner, self-sold products are primarily sourced from Chinese vendors, who typically provide image-based introductions with limited Chinese text embedded in the images. While such image-based content aligns with Chinese consumer preferences, global consumers are more accustomed to text-based descriptive bullet points, such as the “About this item” section on Amazon. Consequently, nearly half of the self-sold products either lack textual descriptions or contain only minimal textual information. During each sub-experiment period, a random subset of consumers who clicked into the product detail pages of the selected products was assigned into the experiment and evenly split to treatment and control groups, resulting in a total of approximately 5 million participants. Control group consumers viewed the original human-generated descriptions, whereas treatment group consumers were shown the GenAI-created descriptions on top of the original, human-created descriptions. [^10] For products without existing human-generated descriptions, control group consumers saw no description—as was historically the case for such products, and treatment group consumers saw the GenAI-generated descriptions only. The treatment group is expected to face lower information asymmetries because GenAI descriptions are more complete, standardized, structured, and accessible.
Marketing Push Message:
@@
The experiment took place over the course of approximately one month in December 2023. A random subset of consumer who received push notifications on their mobile entered into the experiment and were then randomly assigned into either control or treatment groups. Given the large scale of this experiment, we restricted our analysis to the first day, which contained 13 million consumers. On our partner platform, push messages were traditionally created by staff, requiring 1-2 employees several hours each month to produce a few dozen messages. Given the platform’s hundreds of millions of consumers, this limited volume meant that many consumers received identical content, constraining the potential for personalized marketing. Accordingly, in the control group of our experiment, consumers primarily received uniform, human-generated marketing content, totaling roughly 2,000 distinct messages. By contrast, in the treatment group, about 40% of consumers were randomly assigned to GenAI-generated messages, yielding nearly 3 million unique messages and thus far greater differentiation across individuals. The hypothesis in this experiment was that the very large number of GenAI-generated messages would enable the platform to deliver more distinctive marketing content across consumers and achieve refined matching of consumers with messages, thereby leading to better responses.
Google Advertising Title:
@@
The experiment was conducted over twelve days in January 2024, with randomization occurred at the product level. The sample included 3.5 million products selected by the retail platform for advertising in the sponsored section of Google Shopping, representing a diverse set of categories. For Google ads, the quality of the advertisement title is critical: a well-crafted title not only increases product discoverability by aligning with user search keywords but also enhances user clicks by incorporating appealing terms that drive consumer interest. [^11] Many e-commerce platforms maintain libraries of such buzzwords which, based on historical data, are known to boost consumer click-through and conversion rates. In our experiment, the control group retained the original product titles created by the sellers, while in the treatment group, titles used in the ads were optimized by GenAI based on seller titles. [^12] When promoting products on Google Shopping, the platform also used the pricing and image information provided by sellers, and these factors remained unchanged across treatment and control groups in our experiment. A key distinction of this experiment is that the GenAI model was not fine-tuned specifically for the advertising context within the e-commerce domain. As a result, the generated titles may fail to emphasize product attributes most relevant to consumer search and purchase decisions, leading us to adopt a more agnostic stance regarding the expected treatment effect in this setting.
Chargeback Defense:
@@
The experiment, conducted from late October to late December 2023, included over 30 thousand consumers. During this period, a random subset of consumers who initiated chargeback requests was assigned to the experiment and then randomly divided into two groups. Contesting chargeback disputes requires a broad skill set, including claim analysis, evidence collection, and persuasive defense writing, which is especially challenging in cross-border contexts characterized by language barriers and complex regulations and customs. As a result, more than half of chargeback disputes on the focal platform were left unaddressed by sellers. In the control group, consumer claims were initially addressed by sellers. If no action was taken, approximately 3-5 outsourced workers then intervened to resolve the claims. However, these employees could only handle a small fraction of cases elaborately, while most were processed using generalized templates that proved far less effective. In contrast, claims in the treatment group were initially managed by sellers and subsequently supported by GenAI agents. Note that this is the only experiment where costs were not identical between treatment and control conditions. Specifically, in the treatment group, the elimination of 3-5 outsourced employees reduced labor costs, but the effect was negligible. We expect the treatment group to experience higher resolutions, since GenAI can generate more tailored and context-specific responses than template-based staff.
Live Chat Translation:
@@
The experiment was conducted over one month in October 2023 and involved approximately 0.2 million non-English-speaking consumers who contacted the platform’s customer service for issues such as clarifying details of platform-level promotions or resolving disputes with sellers when no agreement was reached. Due to cost constraints, a large share of requests from non-English-speaking consumers are handled by customer service agents from the Philippines who provide assistance in English, as employing native agents for each market is roughly three times more expensive. During the experiment, non-English-speaking consumers who initiated inquiries to the platform’s customer service were randomly split among treatment and control conditions. In the treatment group, consumers interacted with Filipino agents with real-time bidirectional GenAI translation support, while those in the control group engaged with Filipino agents without the aid of translation assistance. We expect the treatment group to face lower communication frictions, as real-time GenAI translation reduces language barriers between consumers and agents, thereby improving service quality and potentially raising conversion rates.
@@
Importantly, in line with the assumptions stated in Section 3.2, labor inputs were held constant between treatment and control conditions when adopting GenAI in each experiment. In particular, there was no or minimal labor displacement across the workflows. Three scenarios account for this outcome. First, in the Pre-sale Service Chatbot and Search Query Refinement experiments, GenAI substituted tasks that were already automated, either through pre-programmed notification or standard search algorithms, and therefore required no labor input. Second, in cases where GenAI augmented rather than replaced existing labor, such as Product Description (where GenAI-generated description was presented ahead of the original human-generated one), Marketing Push Message (where some consumers were randomly assigned to GenAI-generated messages, while others continued receiving human-generated ones), Google Advertising Title (where GenAI optimized ad titles based on human-created product titles), and Live Chat Translation (where agents retained their roles but received real-time GenAI translation support), no reduction in labor inputs occurred. Third, the only exception was the Chargeback Defense experiment, where GenAI adoption technically replaced workers previously responsible for drafting defenses. However, the extent of displacement was negligible, affecting only 3–5 outsourced workers. Moreover, although energy/computing costs may have varied between treatment and control conditions, these were minimal and negligible compared to the cost of other operations. Finally, the adoption of GenAI did not alter product prices.

### v3

@@ 2.1 The Economic Impact of Generative AI
Recent advances in GenAI have attracted considerable attention for their potential economic and social implications. A growing body of research documents GenAI’s ability to enhance individual productivity in simple and well-defined tasks, including mid-level writing [undefaaa], software development [undefaac, undefw], marketing copy generation [undefy], and legal analysis [undefu]. However, this literature largely focuses on individual-level impacts in laboratory settings, with limited empirical evidence of measurable productivity gains at the aggregate or firm level [undef, undefq]. In fact, the productivity dynamics observed in real-world firm adoption are likely more complex than those captured in lab environments due to technical, organizational, and market factors ([undefn]; [undefaw]).
More importantly, most prior studies assess the productivity potential of GenAI from a supply-side perspective, emphasizing labor savings or improvements in worker efficiency, typically measured by decreases in average task completion time or increases in the average number of completed tasks [undefaaa, undefy, undefaac, undefan, e.g.,]. Even the limited studies employing real-world firm data, such as [undefn] and [undefaw], largely rely on similar supply-side metrics. By contrast, very few studies have examined GenAI’s productivity effects through demand-side value creation, such as enhanced consumer experience and increased purchases.777[undeft], [undefad], [undefas], and [undefam] show that AI-generated ad copies or images in digital advertising can raise the number of clicks or click-through rates, but they provide no evidence on actual purchase behavior.
Furthermore, another important question concerns the asymmetric effects of GenAI across different user groups. Earlier waves of technological change, such as the adoption of computers and the Internet, were often described as “skill-biased” ([undefak]), disproportionately favoring skilled users while leaving unskilled workers behind ([undefk]; [undefb]; [undefe]; [undefa]). As of now, the heterogeneous impacts of GenAI on worker performance appear more context-dependent. On one hand, by lowering skill barriers, GenAI can promote inclusivity ([undefay]; [undefac]; [undefv]), as it benefits users with lower levels of skills and expertise ([undefn]; [undefaaa]; [undefaac]; [undefap]). On the other hand, evidence from other studies points to the opposite outcome ([undefaad]; [undefaab]).
@@ 2.2 Friction Reduction in Online Marketplaces
A key friction in online marketplaces is asymmetric information: buyers often cannot directly verify product quality or seller reliability prior to purchase [undefar, undefaai]. To mitigate this challenge and facilitate consumer decision-making, platforms have traditionally relied on reputation and review systems to generate quality signals [undefp, undefab, undefae, undefaak]. However, these feedback mechanisms suffer from well-documented limitations, including grade inflation [undefaz, undefaao], “cold start” problems [undefc], and score manipulation [undefav, undefau]. Modern platforms are increasingly adopting advanced technologies and innovative market designs. For instance, AI models based on natural language processing are used to extract insights from textual reviews or to filter feedback for relevance, enabling more accurate inferences about product quality and consumer satisfaction [undefax, undefat]. Moreover, curated provision of off-site social media information during consumer search has been shown to assist consumer decision-making and enhance platform revenues [undefah].
Consumers on digital platforms also face substantial frictions during online search, arising either from search costs, the effort and resources required to locate information, or search targetability, the effectiveness of search engines in retrieving the most relevant products. Such frictions can significantly influence market outcomes and market structures in online marketplaces [undefai, undefao, undefaal, undefo, undefm, undefd]. To mitigate search frictions, digital platforms have invested heavily in technology and design innovations aimed at encouraging consumer search behavior and improving match value. Prior research highlights several such advances in online search, including ranking algorithms [undefaa, undefaaj, undefaam], refinement tools like sorting and filtering [undefs, undefag], machine-learning-driven personalized search [undefaan], category-refinement-based precision improvements [undefaap], and optimal search engine information layout [undefal].
In online marketplaces, advances in personalization and targeting technologies can also help reduce frictions by delivering products or content to users most likely to be interested. [undefg] develop theoretical models showing that improving advertisers’ targeting ability increases the number of consumer-product matches, thus enhancing the overall social value of advertising. Empirically, targeted advertising has been found more effective than untargeted approaches. For instance, [undefaj] observe that display ads matched to website context significantly raise purchase intent, while [undefh] suggest that targeted ads are particularly valuable in lowering search friction when consumers would otherwise have difficulty discovering or learning about products. Extending these insights to recommendation systems, [undefaag] demonstrate that removing personalized recommendations discourages consumer engagement and purchasing, especially for small sellers and niche consumers.
@@
3 Study Setting, Experimental Design, and Data
@@ 3.2 Theoretical Framework
We model the impact of GenAI adoption on firm productivity through the lens of the standard Solow growth model [undefaae]. Assume that output is produced according to a Cobb–Douglas production function
@@
In Section 3.3, we explain why these assumptions are likely to hold in our context. Under such conditions, the potential gains from GenAI adoption can be interpreted as a pure productivity shock, represented by an upward shift in the $A$ term of the production function. This interpretation is consistent with treatments of past general-purpose technologies (e.g., electrification or the internet), where adoption translated into improvements in TFP rather than capital accumulation. [^8] A natural question is whether GenAI adoption should be viewed as capital deepening rather than TFP growth. While, in principle, new investments in IT infrastructure (e.g., servers, GPUs, or proprietary model development) could expand the capital stock, in our context the platform already possessed the required infrastructure, and GenAI applications constituted incremental software upgrades operating on existing systems. Evidence from consumer-facing sectors further suggests that quality improvements—including reductions in information asymmetry, better matching, enhanced personalization and targeting, and product innovation—can raise revenue-based productivity without corresponding shifts in input units or costs [undefaah, undefl, undefx].
@@ 3.3 Empirical Framework
AI-generated descriptions on top of those created by humans Yes
@@
Pre-sale Service Chatbot
The experiment, conducted over a two-month period from September to October 2023, included a random sample of 44 thousand consumers who initiated pre-sale customer service requests for the platform’s self-sold products during the experimental period. These consumers were randomly divided into treatment and control groups. Consumers in the control group received the platform’s automated response service, which delivered a pre-programmed standardized notification indicating that customer service was unavailable. This auto-response condition reflects the platform’s standard operating practice of prioritizing human agents for post-sale rather than pre-sale support for self-sold products, given that pre-sale inquiries are generally less urgent. This setup also mirrors the constraints faced by many third-party sellers on the platform, particularly smaller-scale sellers, who lack the capacity to provide real-time multilingual customer service. By contrast, consumers in the treatment group were supported purely by GenAI-powered chatbots. GenAI is expected to reduce search costs and asymmetric information between buyers and sellers in the treatment group by providing richer, context-specific responses to consumer inquiries.
Search Query Refinement
@@
The experiment comprised five sub-experiments, each involving consumers who spoke English, Spanish, French, Portuguese, or Korean, which are among the most widely used languages on the platform. All sub-experiments ran for one week in December 2023, with staggered start dates. GenAI was employed to create multilingual, textual product descriptions for a predetermined product set of approximately 45,000 randomly selected platform self-sold products spanning a broad range of categories. [^10] On this platform, product descriptions refer to the text content in the description module of product detail pages that summarizes key features and selling points. According to our partner company, self-sold products are primarily sourced from Chinese vendors, who typically provide image-based product presentations with limited Chinese text embedded in the images. While such image-based content aligns with Chinese consumer preferences, global consumers are more accustomed to text-based descriptive bullet points, such as the “About this item” section on Amazon. Consequently, nearly half of the self-sold products either lack textual descriptions or contain only minimal textual information. During each sub-experiment period, a random subset of consumers who clicked into the product detail pages of the selected products was assigned to the experiment and evenly split into treatment and control groups, resulting in a total of approximately 5 million participants. Control group consumers viewed the original human-generated descriptions, whereas treatment group consumers were shown the GenAI-created descriptions on top of the original, human-created descriptions. [^11] For products without existing human-generated descriptions, control group consumers saw no description—as was historically the case for such products, and treatment group consumers saw the AI-generated descriptions only. The treatment group is expected to face lower information asymmetries, as AI-generated descriptions are more complete, standardized, and structured, and are accessible in multiple languages.
Marketing Push Message
@@
The experiment took place over the course of approximately one month in December 2023. A random subset of consumer who received push notifications on their mobile entered into the experiment and were then randomly assigned into either control or treatment groups. Given the large scale of this experiment, we restricted our analysis to the first day, which contained 13 million consumers. On our partner platform, push messages were traditionally created by staff, requiring 1-2 employees several hours each month to produce a few dozen messages. Given the platform’s hundreds of millions of consumers, this limited volume meant that many consumers received identical content, constraining the potential for personalized marketing. Accordingly, in the control group of our experiment, consumers primarily received uniform, human-generated marketing content, totaling roughly 2,000 distinct messages. By contrast, in the treatment group, about 40% of consumers were randomly assigned to AI-generated messages, yielding nearly 2.7 million unique messages and thus far greater differentiation across individuals. The hypothesis in this experiment was that the very large number of AI-generated messages would enable the platform to deliver more distinctive marketing content across consumers and achieve refined matching of consumers with messages, thereby leading to better responses.
Google Advertising Title
@@
The experiment was conducted over twelve days in January 2024, with randomization occurring at the product level. The sample included 3.5 million products selected by the retail platform for advertising in the sponsored section of Google Shopping, representing a diverse set of categories. For Google ads, the quality of the advertisement title is critical: a well-crafted title not only increases product discoverability by aligning with user search keywords but also enhances user clicks by incorporating appealing terms that drive consumer interest. [^12] Many e-commerce platforms maintain libraries of such buzzwords which, based on historical data, are known to boost consumer click-through and conversion rates. In our experiment, the control group retained the original product titles created by the sellers, while in the treatment group, titles used in the ads were optimized by GenAI based on seller titles. [^13] When promoting products on Google Shopping, the platform also used the pricing and image information provided by sellers, and these factors remained unchanged across treatment and control groups in our experiment. A key distinction of this experiment is that the GenAI model was not fine-tuned specifically for the advertising context within the e-commerce domain. As a result, the generated titles may fail to emphasize product attributes most relevant to consumer search and purchase decisions, leading us to adopt a more agnostic stance regarding the expected treatment effect in this setting.
Chargeback Defense
@@
The experiment, conducted from late October to late December 2023, included over 30 thousand consumers. During this period, a random subset of consumers who initiated chargeback requests was assigned to the experiment and then randomly divided into two groups. Contesting chargeback disputes requires a broad skill set, including claim analysis, evidence collection, and persuasive defense writing, which is especially challenging in cross-border contexts characterized by language barriers and complex regulations and customs. As a result, more than half of chargeback disputes on the focal platform were left unaddressed by sellers. In the control group, consumer claims were initially addressed by sellers. If no action was taken, approximately 3-5 outsourced workers then intervened to resolve the claims. However, these employees could only handle a small fraction of cases elaborately, while most were processed using generalized templates that proved far less effective. In contrast, claims in the treatment group were initially managed by sellers and subsequently supported by GenAI agents. We expect the treatment group to experience higher resolutions, since GenAI can generate more tailored and context-specific responses than template-based staff.
Live Chat Translation
@@
The experiment was conducted over one month in October 2023 and involved approximately 0.2 million non-English-speaking consumers who contacted the platform’s customer service for issues such as clarifying details of platform-level promotions or resolving disputes with sellers when no agreement was reached. Due to cost constraints, a large share of such requests is handled by customer service agents from the Philippines who provide assistance in English, as employing native agents for each market is roughly three times more expensive. During the experiment, non-English-speaking consumers who initiated inquiries to the platform’s customer service were randomly split among treatment and control conditions. In the treatment group, consumers interacted with Filipino agents with real-time bidirectional GenAI translation support, while those in the control group engaged with Filipino agents without GenAI translation assistance. We expect the treatment group to face lower communication frictions, as real-time GenAI translation reduces language barriers between consumers and agents, thereby improving service quality and potentially raising conversion rates.
@@
Importantly, consistent with the assumptions outlined in Section 3.2, labor inputs were held constant across treatment and control conditions in all experiments, and GenAI adoption involved little to no labor displacement, with the exception of one workflow. In the Pre-sale Service Chatbot and Search Query Refinement workflows, GenAI replaced tasks that were already automated, such as pre-programmed notifications or standard search algorithms, and therefore required no labor input. In other workflows, GenAI augmented rather than replaced existing labor. For example, AI-generated content complemented human inputs in Product Description, Marketing Push Message, and Google Advertising Title, while Live Chat Translation provided real-time support without altering agents’ roles. The sole exception is Chargeback Defense, where GenAI replaced workers previously responsible for drafting dispute responses. However, the extent of displacement was negligible, affecting only three to five outsourced workers. Moreover, although energy and computing costs may have differed between treatment and control conditions, these costs were minimal relative to other operating expenses. Finally, GenAI adoption did not alter product prices.